OpenStudy (anonymous):

How would I make a pie chart with this data: putting them in sections 1-5, so here's the data 1500,11000,10000,9000,5000

OpenStudy (anonymous):

(Remember pie charts add up to 100%)

OpenStudy (haleyelizabeth2017):

you might need to divide by 100 or something....... I can try something different....one second

OpenStudy (campbell_st):

just find the total of the data... then its each piece of data divided by total times 100 this gives the percentage then find the percentage of 360 degrees is a revolution. e.g. of the data totals 10000 then 1500/10000 * 100 = 15% 15% of 360 = 54 so 1500 has a sector angle of 54 degrees. 

OpenStudy (anonymous):

This isn't geometry hold on let me send y'all a picture of the problem

OpenStudy (anonymous):

@I sent a pic..... Does this make sense to u @haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

not really.....:/

OpenStudy (anonymous):

I don't blame yah, guess I'll be asking my teacher for help on this one

OpenStudy (here_to_help15):

@haleyelizabeth2017 checks ur websmails

OpenStudy (haleyelizabeth2017):

OMG please stop! My answer WILL NOT change....EVER

OpenStudy (aum):

1000 can be factored out. The data points are: 15, 11, 10, 9, 5 Total = 50 Percentages: 15/50*100 = 15*2 = 30% 11 * 2 = 22% 10 * 2 = 20% 9 * 2 = 18% 5 * 2 = 10%

OpenStudy (haleyelizabeth2017):

good job! I am saying thanks on Blu's behalf :)

OpenStudy (aum):

Enter the following: Section 1 30% Section 2 22% Section 3 20% Section 4 18% Section 5 10%

OpenStudy (anonymous):

Leave ((MY BESTIE ALONEEE OR YOU WILL BE SEEKING REVENGEEE !!!!!!!!!!!!!!!!!!!!!!!))

OpenStudy (anonymous):

thank u aum... I woulda never guessed

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

ok

OpenStudy (haleyelizabeth2017):

Who were you talking to Blu? HTH15?

OpenStudy (anonymous):

That was right, Thank You!

OpenStudy (anonymous):

yes I was referring to him

OpenStudy (haleyelizabeth2017):

:) okay :P